OpenStudy (anonymous):

given f(x) shown below, complete the equation for the inverse of f(x). if necessary use the slash mark (/) for the division problem

OpenStudy (anonymous):

\[f ^{-1}(y)=\]

OpenStudy (anonymous):

f(x)=6x-7

hero (hero):

I showed you how to do these yesterday.

OpenStudy (anonymous):

i got \[f ^{-1}(y)=\frac{ x+6 }{ 7 }\]

hero (hero):

wait

hero (hero):

Okay, that is correct.

OpenStudy (anonymous):

yeah kay ay ill post others can u check them please??????

hero (hero):

Okay.

OpenStudy (anonymous):

given f(x) shown below, complete the equation for the inverse of f(x). if necessary use the slash mark (/) for the division problem F(x)=x-2 \[F ^{-1}(y)=\]

OpenStudy (anonymous):

\[f ^{-1}(y)=x+2\]

OpenStudy (anonymous):

f(x)=8x \[f ^{-1}(y)=\]

OpenStudy (anonymous):

\[f ^{-1}(y)=-8\]

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

yo hero still therr??????????????????????????